A stock is selling for $32 in the stock market, what might the market assuming about its constant growth rate in dividends if the firm is expected to have earnings per share (EPS) of $5 and 40% payout ratio next year. Assume a 13% discount rate
Answer #1

Dividend next year = 5 * 0.4 = $2

Price = Div1 * (1 + g)/(r - g)

32 = 2 * (1 + g)/(0.13 - g)

32 * 0.13 - 32g = 2 + 2g

34g = 2.16

g = 2.16/34

g = 0.06352941176

g = 6.352941176%